# Topological characterization of magnon-polaron bands and thermal Hall conductivity in a frustrated kagome antiferromagnet

### v1

@@ III.1 Bulk spectra and topological transition by magnon-phonon coupling
To visualize this, we present the bulk band structure near the transition points in Fig. 3. Despite finite values of $B_{0}$ and $D$, we observe a gap-closing scenario near the $\mathbf{\Gamma}$ point between the lower and the middle bands along the transition line $\text{u}_{1}$ (see Fig. 2(d)). The bands reopen beyond the $\text{u}_{1}$ contour, as shown in Fig. 3(b), where the Chern number set is given by $C=(1,-2,1)$. Furthermore, with increasing coupling strength $g^{l}$, we observe another gap-closing event at larger values of $g^{l}$ near the $\mathbf{K}$ point in Fig. 3(c), which becomes gapped again beyond the $\text{u}_{2}$ line with the Chern number set $C=(-1,0,1)$, as shown in Fig. 3(d). While we keep the phases corresponding to low values of $J_{2}$ and $D$ (to maintain the $\bm{Q}=0$ LRO) in the main text, the results for relatively larger values of $J_{2}$ have been shown in Appendix B.
@@
Thus far we have investigated multiple phase transitions in the presence of local spin-phonon interaction ($H_{sp}^{l}$), where the involvement of phonon degrees of freedom differs from that in the non-local spin-phonon interaction ($H_{sp}^{nl}$), as given in Eq. (9). Although similar modifications appear in Eq. (12) in the presence of $H_{sp}^{nl}$, the magnitudes of the effective parameters, such as $\lambda_{q}^{nl}$ and $\Delta^{nl}$, differ from those in the local case. Consequently, at $\hbar\omega=J_{1}S$, the interplay between these factors yields distinct outcomes.
@@
To provide further insight into these mechanisms, we also construct the topological phase diagram as a function of $D$ and the non-local spin-phonon coupling strength $g^{nl}$. Here, in Fig. 4(a), we observe two different topological regions with distinct Chern number sets $C=(3,-4,1)$ and $C=(1,-2,1)$. Further, the transition between these regions, marked by the line $\text{l}_{1}$, must involve a gap-closing event. Moreover, prior to crossing $\text{l}_{1}$, a narrow gap separates the lower and middle bands near the $\mathbf{\Gamma}$ point, which becomes gapless along that transition line, as seen from Figs. 4(b) and 4(c). However, beyond the transition line $\text{l}_{1}$, a gap reopens with a distinct Chern number set and the system retains its topological characteristics as shown in Fig. 4(d).
@@
Although in both cases we observe that with increasing $g^{l(nl)}$, the band energy corresponding to the lower band increases near the $\mathbf{\Gamma}$ point. These characteristics depend on the competition between the factors of the renormalized magnon-polaron bands, namely $\Delta^{l(nl)}$ and $e^{-\lambda_{q}^{l(nl)}}$ (Appendix A.2) at a fixed value of the phonon frequency, $\omega$. Therefore, if we increase $\omega$, it may cause the lower band energy to decrease near the $\mathbf{\Gamma}$ point with increasing $g^{l(nl)}$, which may eventually acquire imaginary eigenvalues and induce a magnon-polaron instability near the $\mathbf{\Gamma}$ point above a certain limit of coupling strength. Notably, the key difference in the outcomes for the two distinct couplings is observed in the phase plots, which suggests that tuning the local coupling strength extracts an additional phase (comprising of three distinct phases) compared to the non-local case (probing two distinct phases) for a given range of $g^{l(nl)}$. The plausible reason behind this disparity stems from the difference in the polaronic weights ($\lambda^{l(nl)}$’s) of the magnon-polaron bands carried by these two (see Appendix A.2). While the renormalization by the local coupling becomes more complex and dominant, leading to a strong band-narrowing effect, the non-local coupling involving NN phonon modes gets screened and thus produces weaker effects.
@@ III.2 Chiral magnon-polaron edge modes scenario
So far, we have explored the topological characteristics and phase transitions induced solely by spin-phonon interactions through the analysis of bulk band structures and their associated Chern numbers. However, it is equally important to verify these topological features by examining the edge modes in order to establish the bulk-boundary correspondence. While the magnonic edge modes in kagome antiferromagnets under an external magnetic field, without spin-phonon interactions, have been investigated in literature [20], the present work focuses on hybridized magnon-polaron states rather than the bare magnons. Subsequently, to examine the bulk-boundary correspondence, we consider a two-dimensional semi-infinite ribbon geometry, finite along one direction (for our case, it is the $y$-direction) and infinite along the other ($x$-direction). Further, to analyze the number of crossings between the edge modes as well as their direction of propagation, we compute the winding number ($\mathcal{W}_{n}$), defined as the sum of the Chern numbers up to the $n^{\text{th}}$ band and is given by [100, 26],
Figure 5: The bulk boundary correspondence is illustrated in both types of spin-phonon coupling regimes. In the presence of local spin-phonon interaction, the edge modes are shown for (a) $g^{l}=0.4J_{1}$ and (c) $g^{l}=0.8J_{1}$. Similarly, for the non-local spin-phonon interaction, the edge modes are displayed for (b) $g^{nl}=0.4J_{1}$ and (d) $g^{nl}=0.8J_{1}$. In all the cases, the other parameters are fixed at $B_{0}=0.4B_{s}$, $J_{2}=0.03J_{1}$, and $D=0.045J_{1}$.
@@
Here, to study the fate of the edge states in the topologically gaped regimes of Figs. 3 and 4, we compute $\mathcal{W}_{n}$ for both $H_{sp}^{l}$ and $H_{sp}^{nl}$, using a fixed set of parameters: $B=0.4B_{s}$, $D=0.045J_{1}$, $J_{2}=0.03J_{1}$, and $T=0$, to enable a direct comparison between the two cases. Now, in presence of $H_{sp}^{l}$, with $g^{l}=0.4J_{1}$, the Chern number set associated with the bands is $C=(1,-2,1)$. Therefore, according to Eq. (15), we find $\mathcal{W}_{1}=1$, indicating one crossing of the edge modes in the first spectral gap, with positive chirality, and $\mathcal{W}_{2}=-1$, indicating another such crossing in the second spectral gap, with negative chirality. Furthermore, in Fig. 5(a), due to an overlap between the lower and the middle bands, there is no such crossing corresponding to $\mathcal{W}_{1}$, while another edge crossing corresponding to $\mathcal{W}_{2}$ is clearly visible at $g^{l}=0.4J_{1}$. Further, by increasing $g^{l}$ to $0.8J_{1}$, the system undergoes a topological phase transition, with the Chern number set changing from $C=(1,-2,1)$ to $C=(-1,0,1)$. Hence, applying Eq. (15), we find $\mathcal{W}_{1}=-1$ and $\mathcal{W}_{2}=-1$, indicating a single edge crossing in each of the spectral gaps, where the corresponding edge currents traverse with negative chirality, as shown in Fig. 5(c).
@@
Although the number of edge crossings remains the same in all cases, except for $g^{nl}=0.4J_{1}$, the direction of propagation, determined by the sign of the winding number $\mathcal{W}_{n}$, varies distinctly across different cases considered by us. Furthermore, even with the same spin-phonon coupling strength, the local and non-local scenarios induce different topological phases. It is also noteworthy that in both Figs. 5(a) and 5(b), the lower and the middle bands exhibit overlap along the $k_{x}$, consistent with the bulk band structure, where these bands lie in close proximity of the $\mathbf{\Gamma}$ and the $\mathbf{K}$ points.
@@ III.3 Magnon-polaron thermal Hall conductivity
We have seen so far that in the presence of an external magnetic field, a frustrated kagome antiferromagnet exhibits various topological phase transitions driven by strong spin-phonon coupling involving optical phonons. Although we have characterized the topological states using Chern numbers and the bulk boundary correspondences have been robustly tested, Hall conductivity being one of the measurable quantities in experiments, remains to be explored. Unlike fermionic systems, where the occupied states below the Fermi level contribute to the quantized Hall conductivity, in bosonic systems all the states with positive energy can contribute to the thermal Hall conductivity. Nevertheless, at low temperatures, the dominant contribution arises from the highly occupied low-energy states. The thermal Hall conductivity (scaled by $\frac{\hbar}{k_{B}}$), is given by [101, 13],
@@
where $k_{B}$ is the Boltzmann constant, $T$ is the temperature, and $\rho_{n,\mathbf{k}}=1/[\exp{(\varepsilon_{n}(\mathbf{k})/k_{B}T)}-1]$ denotes the BE distribution function, with $\varepsilon_{n}(\mathbf{k})$ representing the eigenvalue associated with the $n^{\text{th}}$ band. Moreover, $c_{2}(\rho_{n,\mathbf{k}})$ serves as the weighting function, defined as,
$\begin{split}c_{2}(\rho_{n,\mathbf{k}})&=(1+\rho_{n,\mathbf{k}})\left(\log{\frac{1+\rho_{n,\mathbf{k}}}{\rho_{n,\mathbf{k}}}}\right)^{2}-(\log{\rho_{n,\mathbf{k}}})^{2}\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-2\text{Li}_{2}(-\rho_{n,\mathbf{k}}),\end{split}$ (17)
@@
where $Li_{2}(-\rho_{n,\mathbf{k}})$ denotes the polylogarithmic function. Further, the function $c_{2}(\rho_{n,\mathbf{k}})$ attains larger values for lower-energy bands at low temperatures [26]. As a result, the low-energy spectrum dominates the behavior of the thermal Hall conductivity, while contributions from the higher-energy states become insignificant.
@@
As mentioned earlier, the LRO state is stabilized at small values of $J_{2}$ and $D$ relative to $J_{1}$. Moreover, while analyzing the band structures and their topological properties, the external magnetic field ($B_{0}$) has been fixed at $0.4B_{s}$ (equivalent to $2.53J_{1}$), corresponding to a specific classical ground state. However, varying the external magnetic field, while keeping all other parameters fixed at $J_{2}=0.03J_{1}$, $D=0.045J_{1}$ and $\widetilde{T}=0.4$ ($\widetilde{T}=k_{B}T/J_{1}S$), leads to a change in the classical ground state and may give rise to distinct topological phases.
@@
Figure 6: The behavior of the thermal Hall conductivity ($\kappa_{xy}$) in the presence of local spin-phonon interaction is shown. (a) $\kappa_{xy}$ is plotted as a function of external magnetic field for different values of the coupling strength $g^{l}$ at $\widetilde{T}=0.4$. The transition lines $s^{0}$ correspond to $g^{l}=0$, while $s^{1},~{}s^{2}$ denote $g^{l}=0.4$, and $s^{3},~{}s^{4}$ refer to $g^{l}/J_{1}=0.8$. (b) It is also presented as a function of $g^{l}$ at two different temperatures, keeping $B_{0}=0.4B_{s}$, where $v^{1},~{}v^{2}$ denote $\widetilde{T}=0.4$ and $v^{3},~{}v^{4}$ correspond to $\widetilde{T}=0.8$. Additionally, (c) the thermal Hall conductivity is shown as a function of temperature for distinct values of $g^{l}$. Here, $\widetilde{\kappa}_{xy}=\kappa_{xy}/J_{1}S$ and $\widetilde{T}=k_{B}T/J_{1}S$. Throughout this analysis, the other parameters are fixed at $J_{2}=0.03J_{1}$ and $D=0.045J_{1}$.
@@
Further, in Fig. 6(a), we observe that at low values of $B_{0}$, the Hall conductivity is positive, corresponding to the topological state $C=(3,-4,1)$, and increases with increasing $B_{0}$. It reaches a peak near $B_{0}\approx 0.1B_{s}$, where the contribution from the lower band becomes significant. Beyond that, the Hall conductivity begins to decrease in spite of being in the same topological state, indicating a reduction in the spectral gap near the low-energy $\mathbf{\Gamma}$ point between the lower and the middle bands. Furthermore, we observe a topological phase transition near the $s^{0}$ transition line (at $B_{0}=0.42B_{s}$), where the set of Chern numbers changes from $C=(3,-4,1)$ to $C=(1,-2,1)$. This transition is also associated with a reversal in the direction of the Hall current. The thermal Hall conductivity becomes zero at a critical value of $B_{0}$, where the contribution from the Berry curvature (along with that coming from the $c_{2}(\rho_{n,\mathbf{k}})$ function) vanishes.
@@
Now, considering local spin-phonon coupling with $g^{l}=0.4J_{1}$, we observe two topological phase transitions as a function of $B_{0}$, each associated with distinct classical ground states. Moreover, the Hall conductivity also shows discontinuity near that transition line. Prior to the $s^{1}$ ($B_{0}\approx 0.16B_{s}$) transition line, where the Hall conductivity increases as a function of $B_{0}$, the topological state corresponds to $C=(1,0,-1)$, which changes to $C=(1,-2,1)$ beyond the $s^{1}$ transition line. Furthermore, a second phase transition occurs near the $s^{2}$ line (at $B_{0}\approx 0.41B_{s}$), where the topological state changes from $C=(1,-2,1)$ to $C=(-1,0,1)$. Moreover, for a different coupling strength, $g^{l}=0.8J_{1}$, we observe a similar sequence of topological phase transitions involving the same sets of Chern numbers, but occurring at different values of the external magnetic field. Specifically, the first transition occurs near the $s^{3}$ transition line at $B_{0}=0.25B_{s}$, followed by a second transition near the $s^{4}$ line at $B_{0}=0.37B_{s}$. Nevertheless, it is important to mention that, in the presence of spin-phonon coupling, the band topology is no longer independent of temperature. Hence, the same transition points will feature at different temperatures. We shall discuss this in more detail later.
@@
However, based on the band properties and the associated Chern numbers, we have already demonstrated that spin-phonon coupling can induce topological phase transitions in the system. Therefore, in Fig. 6(b), we present the behavior of the Hall conductivity as a function of the coupling strength ($g^{l}$), which reflects the signatures of topological phase transitions in its characteristics. Here, we examine the Hall conductivity as a function of $g^{l}$, varied from 00 to $J_{1}$, at two different temperatures, while keeping the external magnetic field fixed at $B_{0}=0.4B_{s}$. In Fig. 2(d)), we have shown that at $\widetilde{T}=0$, for the same values of $D$ and $J_{2}$ as mentioned earlier, two phase transitions occur as a function of $g^{l}$ within this range. Similarly in Fig. 6(b), at a finite temperature, specifically, at $\widetilde{T}=0.4$, we observe a phase transition from $C=(3,-4,1)$ to $C=(1,-2,1)$ across the $v^{1}$ line at $g^{l}=0.105J_{1}$, whereas the other transition from $C=(1,-2,1)$ to $C=(-1,0,1)$ occurs across the $v^{2}$ line at $g^{l}=0.52J_{1}$. Owing to the temperature dependency on the energy bands, at $\widetilde{T}=0.5$, the transitions occur at different values of $g^{l}$. As $g^{l}$ is varied from 00 to $J_{1}$, the first transition takes place at the $v^{3}$ line for $g^{l}=0.1J_{1}$, while the second transition occurs across the $v^{4}$ line for $g^{l}=0.49J_{1}$. Additionally, with the first transition occurring at a low value of $g^{l}$, the gap closing takes place at the $\mathbf{\Gamma}$ point (see Fig. 3(a)), where the energies corresponding to the lower and the middle bands are very small. Moreover, due to the presence of the $c_{2}(\rho_{n,\mathbf{k}})$ function, the primary contribution in the thermal Hall conductivity comes from that region of the BZ (vicinity of the $\mathbf{\Gamma}$ point), resulting in a noticeable kink near $v^{1(3)}$ transition line. Further, at $v^{2(4)}$, the gap-closing transition occurs at the $\mathbf{K}$ point (see Fig. 3(c)), where the corresponding band energies are relatively high. Hence, the Hall conductivity curve exhibits a comparatively less noticeable deviation near the transition lines. Nevertheless, as the coupling strength $g^{l}$ increases, the Chern number corresponding to the lower band decreases from positive to negative integer values, indicating the emergence of negative Berry curvature of the lower band, which contributes most significantly to the Hall conductivity. Consequently, the thermal Hall conductivity becomes more negative with increasing coupling strength.

### v2

@@ III.1 Bulk spectra and topological transition by spin-phonon coupling
To visualize this, we present the bulk band structure near the transition points in Fig. 4. Despite finite values of $B_{0}$ and $D$, we observe a gap closing scenario near the $\mathbf{\Gamma}$ point between the lower and the middle bands along the transition line $\text{u}_{1}$ (see Fig. 4(a)). The bands reopen beyond the $\text{u}_{1}$ contour, as shown in Fig. 4(b), where the Chern number set is given by $C=(1,-2,1)$. Furthermore, with increasing coupling strength $g^{l}$, we observe another gap closing event at larger values of $g^{l}$ near the $\mathbf{K}$ point in Fig. 4(c), which becomes gapped again beyond the $\text{u}_{2}$ line with the Chern number set $C=(-1,0,1)$, as shown in Fig. 4(d). While we keep the phases corresponding to low values of $J_{2}$ and $D$ (to maintain the $\bm{Q}=0$ LRO phase) in the main text, the results for relatively larger values of $J_{2}$ have been shown in Appendix B.
@@
Thus far we have investigated multiple phase transitions in the presence of local spin-phonon interaction ($H_{\text{sp}}^{l}$), where the involvement of phonon degrees of freedom differs from that in the non-local spin-phonon interaction ($H_{\text{sp}}^{nl}$), as given in Eq. (9). Although similar modifications appear in Eq. (12) in the presence of $H_{\text{sp}}^{nl}$, the magnitudes of the effective parameters, such as $\lambda_{q}^{nl}$ and $\Delta^{nl}$, differ from those in the local case. Consequently, at $\hbar\omega=J_{1}S$, the interplay between these factors yields distinct outcomes.
@@
To provide further insight into these mechanisms, we also construct the topological phase diagram as a function of $D$ and the non-local spin-phonon coupling strength $g^{nl}$. Here, in Fig. 5(a), we observe two different topological regions with distinct Chern number sets $C=(3,-4,1)$ and $C=(1,-2,1)$. Further, the transition between these regions, marked by the line $\text{l}_{1}$, must involve a gap closing event. Moreover, prior to crossing $\text{l}_{1}$, a narrow gap separates the lower and middle bands near the $\mathbf{\Gamma}$ point, which becomes gapless along that transition line, as seen from Figs. 5(b) and 5(c). However, beyond the transition line $\text{l}_{1}$, a gap reopens with a distinct Chern number set and the system retains its topological characteristics as shown in Fig. 5(d).
@@
Although in both cases we observe that with increasing $g^{l(nl)}$, the band energy corresponding to the lower band increases near the $\mathbf{\Gamma}$ point. These characteristics depend on the competition between the factors of the renormalized magnon-polaron bands, namely $\Delta^{l(nl)}$ and $e^{-\lambda_{q}^{l(nl)}}$ (Appendix A.2) at a fixed value of the phonon frequency, $\omega$. Therefore, if we increase $\omega$, it may cause the lower band energy to decrease near the $\mathbf{\Gamma}$ point with increasing $g^{l(nl)}$, which may eventually acquire imaginary eigenvalues and induce a magnon-polaron instability near the $\mathbf{\Gamma}$ point above a certain limit of coupling strength. Notably, the key difference in the outcome for the two distinct couplings is observed from the phase plots, which indicates that tuning of the local coupling strength ($g^{l}$) yields an additional phase (phase diagram comprises of three distinct phases) compared to the non-local case ($g^{nl}$), which renders two distinct phases. The plausible reason behind this disparity stems from the difference in the polaronic weights ($\lambda^{l(nl)}_{q}$) of the magnon-polaron bands corresponding to these two cases (see Appendix A.2). While the renormalization due to the local coupling gives rise to more complex and dominant effects, leading to strong band-narrowing features, the non-local coupling involving NN phonon modes suffers screening and thus produces weaker effects.
@@ III.2 Chiral magnon-polaron edge modes scenario
So far, we have explored the topological characteristics and phase transitions induced solely by spin-phonon interactions through the analysis of bulk band structures and their associated Chern numbers. However, it is equally important to verify these topological features by examining the edge modes in order to establish the bulk-boundary correspondence. While the magnonic edge modes in kagome antiferromagnets under an external magnetic field, without spin-phonon interactions, have been investigated in literature [20], the present work focuses on hybridized magnon-polaron states rather than the bare magnons. Subsequently, to examine the bulk-boundary correspondence, we consider a two-dimensional semi-infinite ribbon geometry, finite along one direction (for our case, it is the $y$-direction) and infinite along the other ($x$-direction). Further, to analyze the number of crossings between the edge modes as well as their direction of propagation, we compute the winding number ($\mathcal{W}_{n}$), defined as the sum of the Chern numbers up to the $n^{\text{th}}$ band and is given by [100, 26]
Figure 6: The edge states of the magnon-polaron bands corresponding to both types of spin-phonon interaction are shown. In the presence of local spin-phonon coupling, the edge modes are plotted for (a) $g^{l}=0.4J_{1}$, and (c) $g^{l}=0.8J_{1}$. Similarly, for the non-local spin-phonon coupling, the edge modes are displayed for (b) $g^{nl}=0.4J_{1}$ and (d) $g^{nl}=0.8J_{1}$. In all the cases, the other parameters are fixed at $B_{0}=0.4B_{s}$, $J_{2}=0.03J_{1}$, and $D=0.045J_{1}$ at zero temperature. These figures establish the bulk-boundary correspondence for our study.
@@
Here, to study the fate of the edge states in the topologically gapped regimes of Figs. 4 and 5, we compute $\mathcal{W}_{n}$ for both $H_{\text{sp}}^{l}$ and $H_{\text{sp}}^{nl}$, using a fixed set of parameters: $B=0.4B_{s}$, $D=0.045J_{1}$, $J_{2}=0.03J_{1}$, and $T=0$, to enable a direct comparison between the two cases. Now, in presence of $H_{\text{sp}}^{l}$, with $g^{l}=0.4J_{1}$, the Chern number set associated with the bands is $C=(1,-2,1)$. Therefore, according to Eq. (15), we find $\mathcal{W}_{1}=1$, indicating one crossing of the edge modes in the first spectral gap, with positive chirality, and $\mathcal{W}_{2}=-1$, indicating another such crossing in the second spectral gap, with negative chirality. Furthermore, in Fig. 6(a), due to an overlap between the lower and the middle bands, there is no such crossing corresponding to $\mathcal{W}_{1}$, while another edge crossing corresponding to $\mathcal{W}_{2}$ is clearly visible at $g^{l}=0.4J_{1}$. Further, by increasing $g^{l}$ to $0.8J_{1}$, the system undergoes a topological phase transition, with the Chern number set changing from $C=(1,-2,1)$ to $C=(-1,0,1)$. Hence, applying Eq. (15), we find $\mathcal{W}_{1}=-1$ and $\mathcal{W}_{2}=-1$, indicating a single edge crossing in each of the spectral gaps, where the corresponding edge currents traverse with negative chirality, as shown in Fig. 6(c).
@@
Although the number of edge crossings remains the same in all cases, except for $g^{nl}=0.4J_{1}$, the direction of propagation, determined by the sign of the winding number $\mathcal{W}_{n}$, varies distinctly across different cases considered by us. Furthermore, even with the same spin-phonon coupling strength, the local and non-local scenarios induce different topological phases. It is also noteworthy that in both Figs. 6(a) and 6(b), the lower and the middle bands exhibit overlap along the $k_{x}$, consistent with the bulk band structure, where these bands lie in close proximity of the $\mathbf{\Gamma}$ and the $\mathbf{K}$ points.
@@ III.3 Magnon-polaron thermal Hall conductivity
We have seen so far that in the presence of an external magnetic field, a frustrated kagome antiferromagnet exhibits various topological phase transitions driven by strong spin-phonon coupling involving optical phonons. Although we have characterized the topological states using Chern numbers and the bulk-boundary correspondences have been robustly tested, Hall conductivity, being one of the measurable quantities in experiments, remains to be explored. Unlike fermionic systems, where the occupied states below the Fermi level contribute to the quantized Hall conductivity, in bosonic systems, all the states with positive energy can contribute to the thermal Hall conductivity. Nevertheless, at low temperatures, the dominant contribution arises from the highly occupied low-energy states. The thermal Hall conductivity (scaled by $\frac{\hbar}{k_{B}}$) is given by [101, 13]
@@
where $k_{B}$ is the Boltzmann constant, $T$ is the temperature, and $\rho_{n,\mathbf{k}}=1/[\exp{(\varepsilon_{n}(\mathbf{k})/k_{B}T)}-1]$ denotes the BE distribution function, with $\varepsilon_{n}(\mathbf{k})$ representing the eigenvalue associated with the $n^{\text{th}}$ band. Moreover, $c_{2}(\rho_{n,\mathbf{k}})$ serves as the weighting function, defined as
$c_{2}(\rho_{n,\mathbf{k}})$ $=$ $(1+\rho_{n,\mathbf{k}})\left(\log{\frac{1+\rho_{n,\mathbf{k}}}{\rho_{n,\mathbf{k}}}}\right)^{2}-(\log{\rho_{n,\mathbf{k}}})^{2}$ (17)
$-2\text{Li}_{2}(-\rho_{n,\mathbf{k}}),$
@@
where $\text{Li}_{2}(-\rho_{n,\mathbf{k}})$ denotes the polylogarithmic function. Further, the function $c_{2}(\rho_{n,\mathbf{k}})$ attains larger values for lower-energy bands at low temperatures [26]. As a result, the low-energy spectrum dominates the behavior of the thermal Hall conductivity, while contributions from the higher-energy states become insignificant.
@@
As mentioned earlier, the LRO phase can be stabilized at small values of $J_{2}$ and $D$ relative to $J_{1}$. Moreover, while analyzing the band structures and their topological properties, the external magnetic field ($B_{0}$) has been fixed at $0.4B_{s}$ (equivalent to $2.53J_{1}$), corresponding to a specific classical ground state. However, varying the external magnetic field, while keeping all other parameters fixed at $J_{2}=0.03J_{1}$, $D=0.045J_{1}$, $\widetilde{T}=0.4$ ($\widetilde{T}=k_{B}T/J_{1}S$), and $\hbar\omega=J_{1}S$ [102], leads to a change in the classical ground state and may give rise to distinct topological phases.
@@
Figure 7: The effects of local spin-phonon interaction on the behavior of the thermal Hall conductivity ($\kappa_{xy}$) are presented. (a) $\kappa_{xy}$ is plotted as a function of external magnetic field for different values of the coupling strength $g^{l}$ at $\widetilde{T}=0.4$. The transition lines $s^{0}$ correspond to $g^{l}=0$, while $s^{1},~{}s^{2}$ denote $g^{l}/J_{1}=0.4$, and $s^{3},~{}s^{4}$ refer to $g^{l}/J_{1}=0.8$. (b) The variations of $\kappa_{xy}$ are shown as a function of $g^{l}$ at two different temperatures, keeping $B_{0}=0.4B_{s}$, where $v^{1},~{}v^{2}$ denote $\widetilde{T}=0.4$ and $v^{3},~{}v^{4}$ correspond to $\widetilde{T}=0.8$. (c) The same is also demonstrated as a function of temperature for distinct values of $g^{l}$. Here, $\widetilde{\kappa}_{xy}=\kappa_{xy}/J_{1}S$ and $\widetilde{T}=k_{B}T/J_{1}S$. Throughout this analysis, the other parameters are fixed at $J_{2}=0.03J_{1}$, $D=0.045J_{1}$ and $\hbar\omega=J_{1}S$.
@@
In Fig. 7(a), we observe that at low values of $B_{0}$, the Hall conductivity is positive, corresponding to the topological state $C=(3,-4,1)$, and increases with increasing $B_{0}$. It reaches a peak near $B_{0}\approx 0.1B_{s}$, where the contribution from the lower band becomes significant. Beyond that, the Hall conductivity begins to decrease in spite of being in the same topological state, indicating a reduction in the spectral gap near the low-energy $\mathbf{\Gamma}$ point between the lower and the middle bands. Furthermore, we observe a topological phase transition near the $s^{0}$ transition line (at $B_{0}=0.42B_{s}$), where the set of Chern numbers changes from $C=(3,-4,1)$ to $C=(1,-2,1)$. This transition is also associated with a reversal in the direction of the Hall current. The thermal Hall conductivity becomes zero at a critical value of $B_{0}$, where the contribution from the Berry curvature (along with that coming from the $c_{2}(\rho_{n,\mathbf{k}})$ function) vanishes.
@@
Now, considering local spin-phonon coupling with $g^{l}=0.4J_{1}$, we observe two topological phase transitions as a function of $B_{0}$, each associated with distinct classical ground states. Moreover, the Hall conductivity also shows discontinuity near that transition line. Prior to the $s^{1}$ ($B_{0}\approx 0.16B_{s}$) transition line, where the Hall conductivity increases as a function of $B_{0}$, the topological state corresponds to $C=(1,0,-1)$, which changes to $C=(1,-2,1)$ beyond the $s^{1}$ transition line. Furthermore, a second phase transition occurs near the $s^{2}$ line (at $B_{0}\approx 0.41B_{s}$), where the topological state changes from $C=(1,-2,1)$ to $C=(-1,0,1)$. Moreover, for a different coupling strength, $g^{l}=0.8J_{1}$, we observe a similar sequence of topological phase transitions involving the same sets of Chern numbers, but occurring at different values of the external magnetic field. Specifically, the first transition occurs near the $s^{3}$ transition line at $B_{0}=0.25B_{s}$, followed by a second transition near the $s^{4}$ line at $B_{0}=0.37B_{s}$. Nevertheless, it is important to mention that, in the presence of spin-phonon coupling, the band topology is no longer independent of temperature. Hence, the same transition points will not feature at different temperatures. We shall discuss this in more detail later.
@@
However, based on the band properties and the associated Chern numbers, we have already demonstrated that spin-phonon coupling can induce topological phase transitions in the system. Therefore, in Fig. 7(b), we present the behavior of the Hall conductivity as a function of the coupling strength ($g^{l}$), which reflects the signatures of topological phase transitions in its characteristics. Here, we examine the Hall conductivity as a function of $g^{l}$, varied from 00 to $J_{1}$, at two different temperatures, while keeping the external magnetic field fixed at $B_{0}=0.4B_{s}$. In Fig. 3(d), we show that at $\widetilde{T}=0$, for the same values of $D$ and $J_{2}$ as mentioned earlier, two phase transitions occur as a function of $g^{l}$ within this range. Similarly in Fig. 7(b), at a finite temperature, specifically, at $\widetilde{T}=0.4$, we observe a phase transition from $C=(3,-4,1)$ to $C=(1,-2,1)$ across the $v^{1}$ line at $g^{l}=0.105J_{1}$, whereas the other transition from $C=(1,-2,1)$ to $C=(-1,0,1)$ occurs across the $v^{2}$ line at $g^{l}=0.52J_{1}$. Owing to the temperature dependency on the energy bands, at $\widetilde{T}=0.5$, the transitions occur at different values of $g^{l}$. As $g^{l}$ is varied from 00 to $J_{1}$, the first transition takes place at the $v^{3}$ line for $g^{l}=0.1J_{1}$, while the second transition occurs across the $v^{4}$ line for $g^{l}=0.49J_{1}$. Additionally, with the first transition occurring at a low value of $g^{l}$, the gap closing takes place at the $\mathbf{\Gamma}$ point (see Fig. 4(a)), where the energies corresponding to the lower and the middle bands are very small. Moreover, due to the presence of the $c_{2}(\rho_{n,\mathbf{k}})$ function, the primary contribution in the thermal Hall conductivity comes from that region of the BZ (vicinity of the $\mathbf{\Gamma}$ point), resulting in a noticeable kink near $v^{1(3)}$ transition line. Further, at $v^{2(4)}$, the gap closing transition occurs at the $\mathbf{K}$ point (see Fig. 4(c)), where the corresponding band energies are relatively high. Hence, the Hall conductivity curve exhibits a comparatively less noticeable deviation near the transition lines. Nevertheless, as the coupling strength $g^{l}$ increases, the Chern number corresponding to the lower band decreases from positive to negative integer values, indicating the emergence of negative Berry curvature of the lower band, which contributes most significantly to the Hall conductivity. Consequently, the thermal Hall conductivity becomes more negative with increasing coupling strength.